# Quasiconservation Laws and Suppressed Transport in Weakly Interacting Localized Models

### v1

@@ III Quasi-LIOMs
One important note is that one could, in principle, analyze higher $l-$order LIOM corrections $\tilde{n}^{(l)}_{\alpha}$, such that the quasiconservation law holds up to $O(\lambda^{l})$ for some $l\in\mathbb{Z^{+}}$. These corrections will be $(2l+2)$–fermion terms, and their numerical calculations will scale at least as $L^{2l+2}$, becoming very expensive. Furthermore, the energy resonances that arise from perturbation theory will grow increasingly complex. As a result, we do not attempt higher-order calculations in this work.
@@ III.2 Methods
For the disordered model, calculations are presented with a distribution over independent and randomly generated disorder realizations. We also obtain distributions for the quasiperiodic model. Although the model is deterministic, our calculations found that changing the phase shift $\delta$ affects the results. To avoid commensuration issues, we found it useful to sample over these variations with a distribution over the phase $\delta$, over 500 evenly spaced phases between $0<\delta<1/k$, which covers one period of the potential. This setup also allows subsequent analysis of interaction corrections to charge transport in both models.
@@ III.3.1 Disordered Model AGP
Figure 1: Analysis of the AGP Frobenius norm in the interacting Anderson Model. All three figures were generated with 500 disorder realizations and open boundary conditions. (a) Calculated probability distributions of $\ln\|\hat{X}\|_{F}$ with disorder strength $W=2.1$. Each colored curve represents varying system sizes $L$, from $L=20$ to $L=100$ in steps of $10$. (b) Same data as in (a) but with the vertical scale shown on a log scale. The red line represents a polynomial $\|\hat{X}\|_{F}^{-2}$ fit to the tail of the probability distribution of $\|\hat{X}\|_{F}$—see text around Eq. (21) for details on the form plotted. (c) Median of the AGP norm as a function of system size $L$. Each colored curve represents a different disorder strength $W$, from $W=0.1$ to $W=3.1$ in step sizes of 0.5. Here, dashed lines represent power-law fits to the data, $\|\hat{X}\|_{F}=AL^{D}+C$ with power-law exponents $D$. In order of increasing disorder strength, we find $D=2.33,2.33,1.99,1.71,1.45,1.38,1.34$, consistent with expected eventual linear in $L$ scaling for a strictly localized system (see text in Sec. III.3.1 for discussion). The error bars on the median curves, estimated using a bootstrapping method, are very small and are therefore omitted in this and subsequent plots for clarity.
@@ III.3.2 Quasiperiodic Model AGP
Figure 2: Analysis of the AGP Frobenius norm in the interacting quasiperiodic model. All three figures were generated with 500 uniformly sampled phase shifts $\delta$ and open boundary conditions. We refer the reader to Sec. III.2 for details on sampling for the quasiperiodic model. (a) Calculated probability distributions of $\ln\|\hat{X}\|_{F}$ with potential depth $h=3.1$. Each colored distribution represents varying system sizes $L$, from $L=20$ to $L=100$, in steps of 10. (b) Same data as in (a) but with the vertical scale shown on a log scale. The red curve represents a polynomial $\|\hat{X}\|_{F}^{-1.75}$ fit to the tail of the probability distribution of $\|\hat{X}\|_{F}$—see text around Eq. (21) for details on the form plotted. (c) Median of $\|\hat{X}\|_{F}$ as a function of system size $L$. Each colored curve represents a different potential depth $h$, from $h=0.1$ to $h=3.1$ in steps of 0.5. Here, we observe polynomial growth with $L$ of $\mathrm{med}\|\hat{X}\|_{F}$ and an anomalous peak at $L=40$ in the regime $h<2$. For a full discussion of these results, see Sec. III.3.2.
@@
When the states are delocalized at low disorders for $h<2$, one also observes polynomial growth of $\|\hat{X}\|_{F}$. Curiously, in this delocalized regime, $\mathrm{med}\|\hat{X}\|_{F}$ grows non-monotonically, and $\mathrm{med}\|\hat{X}\|_{F}$ spikes at certain system sizes. In the case with $k=\sqrt{2}$, this occurs at $L=40$. These spikes shift to different system sizes when the wavenumber is changed to different irrational numbers, which we have separately checked for $k=\phi$, $k=\sqrt{3}$, and $k=\sqrt{6}$. Here, $\phi$ is the golden ratio. Anomalous behavior of this kind is known to arise in the quasiperiodic model at certain integer distances. For example, Ref. [37] observed that single-particle resonances are enhanced at orbitals separated by distances equal to the Fibonacci numbers $F_{n}$, when one picks $k=\phi$. Similarly, Ref. [74] reported suppressed interaction between LIOMs separated at the Fibonacci numbers for $k=\phi$, and the corresponding integer sequences that relate to the silver and bronze ratios. In our case, we find a more general condition that explains these anomalous system sizes, encompassing previous observations: they occur when the underlying potential is nearly commensurate with the lattice. For open boundary conditions, this corresponds to when $k(L+1)\approx n$ for some $n\in\mathbb{Z}$, so $k\approx\frac{n}{L+1}$ for some integer $n$. For example, in the calculation presented in Fig. 2(c), which was done for $k=\sqrt{2}$, there is a notable spike at $L=40$, and $\sqrt{2}\times 41\approx 58$. This observation helps explain why corresponding integer sequences arise for a given irrational number $k$, an example being the Fibonacci numbers when one chooses $k=\phi$, the golden ratio. Here, we note that $41$ is a member of the integer sequence $\frac{1}{2}Q_{n}$, where $Q_{n}$ are the Pell-Lucas numbers. Just as successive ratios of the Fibonacci numbers form the best rational approximations for $\phi$ ($\lim_{n\rightarrow\infty}F_{n+1}/F_{n}=\phi$), similarly, $\lim_{n\rightarrow\infty}\frac{1}{2}Q_{n}/P_{n}=\sqrt{2}$, where $P_{n}$ here refers to the Pell numbers. Thus, if $L+1=\frac{1}{2}Q_{n}$ for some $n$, then $(L+1)\times\sqrt{2}\approx m\in\mathbb{Z}$, leading to the potential being nearly commensurate with the lattice. We confirm that for $k=\phi$, these spikes occur at system sizes corresponding to the Fibonacci numbers and other integers $m$ that happen to satisfy $m\times k\approx n$. This underlying commensurability is likely the culprit that leads to increased resonances and other strange behavior in the spectrum, leading to a spike in the quantities calculated here. This would be expected as a periodic system exhibits vastly different physics from an incommensurate one.
@@ III.4.1 LIOM Corrections: Anderson Model
Figure 3: Analysis of the Frobenius norm of the first-order corrections to the noninteracting LIOMs, $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$, for the weakly-interacting Anderson model. Each plot includes all the orbitals $\alpha=1,\dots,L$ for each of $500$ disorder realizations. (a)-(d): Distributions of $\ln(\|\tilde{n}_{\alpha}^{(1)}\|_{F})$ at $W=0.6,1.6,2.6,3.1$ and for $\alpha=1,\dots,L$. Within each panel, each colored curve represents a different system size. We show calculations between $L=20$ and $L=100$ in steps of $10$. The inset in (d) shows the same data as in the main (d) panel, but with the vertical scale shown on log-scale. The red line in the inset represents a polynomial $\|\tilde{n}_{\alpha}^{(1)}\|_{F}^{-2}$ fit to the tail of the probability distribution of $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$— see the text surrounding Eq. (21) for details on the form plotted. The convergence of the distributions for large enough $W$ and $L$ indicates the presence of an extensive number of quasiconservation laws after interactions are added, even in the thermodynamic limit. (e): Median $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ as a function of $L$ on a log-log scale. Each colored curve represents a different disorder strength $W$, from $W=0.1$ to $W=3.1$ in steps of $0.5$. One observes that the curves appear to saturate for large $L$, indicating that at least half of the LIOMs are conserved to second order when interactions are added. See Sec. III.4.1 for a full discussion of the results.
We present numerical calculations of the Frobenius norm of the LIOM corrections, $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$, in Fig. 3, for the weakly–interacting Anderson model. For each $L$, we sample $500$ disorder realizations and record $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ for every orbital $\alpha=1,2,...,L$ in each disorder realization, giving a total of $500\times L$ datapoints in each distribution.
@@
Furthermore, in Fig. 4, we note that the $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ vs $L$ curves collapse when the x-axis is appropriately scaled by the participation ratio of the Anderson model, which scales approximately as $\mathrm{PR}(W)\propto W^{-1.78}$ in the range of $W$ considered here (see Appendix E for a more detailed explanation) and when the y-axis is scaled phenomenologically with $\sim\mathrm{PR}^{1.7}$. Here, the participation ratio is used as a proxy for the localization length $\xi$. The collapse of these curves suggests that at low disorders, the growth of $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ is a finite-size effect. If the system size $L$ exceeds the localization length, for any disorder strength $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ is expected to saturate to a finite value. Put together, these results suggest that even in the thermodynamic limit, a fraction of noninteracting LIOMs have finite and small corrections in the presence of interactions.
@@ III.4.2 LIOM Corrections: Quasiperiodic Model
Figure 5: Analysis of the Frobenius norm of the first-order corrections to the noninteracting LIOMs, $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$, for the weakly-interacting quasiperiodic model. (a)-(d): Distributions of $\ln(\|\tilde{n}_{\alpha}^{(1)}\|_{F})$ over 500 phase shifts $\delta$, sampled at potential depths $h=0.6,1.6,2.6,3.1$, with $\alpha=1\dots L$. Each curve represents the distribution calculated at different system sizes, from $L=20$ to $L=100$ in steps of 10. The inset in (d) shows the same data as in the main (d) panel but with the vertical scale shown on a log scale. The red line in the inset represents a polynomial $\|\tilde{n}_{\alpha}^{(1)}\|_{F}^{-1.75}$ fit to the tail of the probability distribution of $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$— see the text surrounding Eq. (21) for details on the form plotted. Similarly to what was observed for the interacting Anderson model, the quasiperiodic model seems to admit an extensive number of quasiconservation laws for the localized phase $h>2$. (e): Median $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ as a function of $L$ on a log-log scale. Each colored curve represents a different potential depth $h$, from $h=0.1$ to $h=3.1$ in steps of 0.5. One observes that the curves for $h<2$ grow with $L$, indicating that the corrections are, unsurprisingly, delocalized. For $h>2$ the curves saturate, indicating that quasi-LIOMs persist when interactions are added to the localized regime of the quasiperiodic model. For a full discussion of the results, see Sec. III.4.2
@@ IV.1 Charge Transport Capacity
In this section, we introduce a measure of the charge transport capacity (for any process) across a link, inspired by the integrated energy current operator introduced in Ref. [75]. We consider an OBC chain and fix a site $i_{0}$ where $11\leq i_{0}\leq L-$ 1. Here, we fix $t=1$. We can define the charge current operator $\hat{J}_{i_{0}}$ across the link between $i_{0}$ and $i_{0}+1$, with rightward current being defined as positive:
@@
Continuing with general discussion, we can now more formally argue that in the MBL phase, the above $\hat{M}_{i_{0}}$ is related to LIOMs and noncommuting partners (raising/lowering operators) localized near $i_{0}$ and is localized near $i_{0}$, and hence the charge transport capacity is bounded by a fixed number in the thermodynamic limit. Indeed, let us take the LIOM description posited in Eq. (7) and express the total charge operator $\hat{N}_{\text{tot}}=\sum_{j}\hat{n}_{j}$ in terms of $\tilde{d}_{\alpha},\tilde{d}_{\alpha}^{\dagger}$. Here, $\tilde{d}_{\alpha},\tilde{d}_{\alpha}^{\dagger}$ are the annihilation and creation operators associated with the corresponding LIOM $\tilde{n}_{\alpha}$. Since $\hat{N}_{\text{tot}}$ commutes with the Hamiltonian, and assuming the Hamiltonian spectrum is non-degenerate, it follows that $\hat{N}_{\text{tot}}$ can be expressed as a sum of terms containing only $\tilde{n}_{\alpha}$’s. [^3] If we could in addition assume that there is an “adiabatic-type” one-to-one correspondence between the unperturbed LIOMs $\hat{d}_{\alpha}^{\dagger}\hat{d}_{\alpha}$ and the new LIOMs $\tilde{d}_{\alpha}^{\dagger}\tilde{d}_{\alpha}$, more specifically that there is a quasi-local unitary $U$ relating $\tilde{n}_{\alpha}\equiv\tilde{d}_{\alpha}^{\dagger}\tilde{d}_{\alpha}=U\hat{d}_{\alpha}^{\dagger}\hat{d}_{\alpha}U^{\dagger}$, requiring also that $U$ commutes with $\hat{N}_{\text{tot}}$, we would conclude $\hat{N}_{\text{tot}}=\sum_{\alpha}\tilde{n}_{\alpha}$. However, this specific form is not used in the argument that follows.
Consider now the parts $\hat{N}_{L}(i_{0})$ and $\hat{N}_{R}(i_{0})$, $\hat{N}_{L}(i_{0})+\hat{N}_{R}(i_{0})=\hat{N}_{\text{tot}}$. The expected quasi-locality of the LIOMs implies that any local operator—in particular, each $\hat{n}_{j}$—is expressible in terms of quasi-local strings containing products of $\tilde{d}_{\alpha},\tilde{d}_{\alpha}^{\dagger},\tilde{n}_{\alpha}$ localized near $j$. For simplicity, let us assume that there is a maximal range $r_{\text{max}}$ for these strings. Now recall that $\hat{N}_{\text{tot}}$ contains only strings of $\tilde{n}$’s. Since any string in $\hat{N}_{R}(i_{0})$ that starts farther than $r_{\text{max}}$ away from $i_{0}$ is not present in $\hat{N}_{L}(i_{0})$ and hence must be the same as in $\hat{N}_{\text{tot}}$, such a string contains only $\tilde{n}$’s and hence commutes with $\hat{H}$. Hence $\hat{M}_{i_{0}}$, which is the off-diagonal part of $\hat{N}_{R}(i_{0})$, obtains contributions only from the strings in $\hat{N}_{R}(i_{0})$ that are localized near $i_{0}$, so $\hat{M}_{i_{0}}$ is indeed localized near $i_{0}$. See Fig. 7 for a visual depiction of this argument. We note that the restriction to strictly local strings (with size at most $r_{\text{max}}$) vs quasi-local (whose amplitudes decrease sufficiently fast with the string size) is not important for this argument.

### v2

@@ III Quasi-LIOMs
[…]
@@ III.2 Methods
For the disordered model, calculations are presented with a distribution over independent and randomly generated disorder realizations. We also obtain distributions for the quasiperiodic model. Although the model is deterministic, our calculations found that changing the phase shift $\delta$ affects the results. To avoid commensuration issues, we found it useful to sample over these variations with a distribution over the phase $\delta$, spanning 500 evenly spaced phases between $0<\delta<1/k$, which covers one period of the potential. This setup also allows subsequent analysis of interaction corrections to charge transport in both models.
@@ III.3.1 Disordered Model AGP
Figure 1: Analysis of the AGP Frobenius norm in the interacting Anderson Model. All three figures were generated based on 500 disorder realizations and open boundary conditions. (a) Calculated probability distributions of $\ln\|\hat{X}\|_{F}$ with disorder strength $W=2.1$. Each colored curve represents varying system sizes $L$, from $L=20$ to $L=100$ in steps of 101010. (b) Same data as in (a) but with the vertical scale shown on a log scale. The red line represents a polynomial $\|\hat{X}\|_{F}^{-2}$ fit to the tail of the probability distribution of $\|\hat{X}\|_{F}$—see text around Eq. (21) for details on the form plotted. (c) Median of the AGP norm as a function of system size $L$. Each colored curve represents a different disorder strength $W$, from $W=0.1$ to $W=3.1$ in step sizes of 0.5. Here, dashed lines represent power-law fits to the data, $\|\hat{X}\|_{F}=AL^{D}+C$ with power-law exponents $D$. In order of increasing disorder strength, we find $D=2.33,2.33,1.99,1.71,1.45,1.38,1.34$, consistent with expected eventual linear in $L$ scaling for a strictly localized system (see text in Sec. III.3.1 for discussion). The error bars on the median curves, estimated using a bootstrapping method, are very small and are therefore omitted in this and subsequent plots for clarity.
@@ III.3.2 Quasiperiodic Model AGP
Figure 2: Analysis of the AGP Frobenius norm in the interacting quasiperiodic model. All three figures were generated based on 500 uniformly sampled phase shifts $\delta$ and open boundary conditions. We refer the reader to Sec. III.2 for details on sampling for the quasiperiodic model. (a) Calculated probability distributions of $\ln\|\hat{X}\|_{F}$ with potential depth $h=3.1$. Each colored distribution represents varying system sizes $L$, from $L=20$ to $L=100$, in steps of 10. (b) Same data as in (a) but with the vertical scale shown on a log scale. The red curve represents a polynomial $\|\hat{X}\|_{F}^{-1.75}$ fit to the tail of the probability distribution of $\|\hat{X}\|_{F}$—see text around Eq. (21) for details on the form plotted. (c) Median of $\|\hat{X}\|_{F}$ as a function of system size $L$. Each colored curve represents a different potential depth $h$, from $h=0.1$ to $h=3.1$ in steps of 0.5. Here, we observe polynomial growth with $L$ of $\mathrm{med}\|\hat{X}\|_{F}$ and an anomalous peak at $L=40$ in the regime $h<2$. For a full discussion of these results, see Sec. III.3.2.
@@
When the states are delocalized at low disorders for $h<2$, one also observes polynomial growth of $\|\hat{X}\|_{F}$. Curiously, in this delocalized regime, $\mathrm{med}\|\hat{X}\|_{F}$ grows non-monotonically, and $\mathrm{med}\|\hat{X}\|_{F}$ spikes at certain system sizes. In the case with $k=\sqrt{2}$, this occurs at $L=40$. These spikes shift to different system sizes when the wavenumber is changed to different irrational numbers, which we have separately checked for $k=\phi$, $k=\sqrt{3}$, and $k=\sqrt{6}$. Here, $\phi$ is the golden ratio. Anomalous behavior of this kind is known to arise in the quasiperiodic model at certain integer distances. For example, Ref. [37] observed that single-particle resonances are enhanced at orbitals separated by distances equal to the Fibonacci numbers $F_{n}$, when one picks $k=\phi$. Similarly, Ref. [74] reported suppressed interaction between LIOMs separated at the Fibonacci numbers for $k=\phi$, and the corresponding integer sequences that relate to the silver and bronze ratios. In our case, we find a more general condition that explains these anomalous system sizes, encompassing previous observations: they occur when the underlying potential is nearly commensurate with the lattice. For open boundary conditions, this corresponds to when $k(L+1)\approx n$ for some $n\in\mathbb{Z}$, so $k\approx\frac{n}{L+1}$ for some integer $n$. For example, in the calculation presented in Fig. 2(c), which was done for $k=\sqrt{2}$, there is a notable spike at $L=40$, and $\sqrt{2}\times 41\approx 58$. This observation helps explain why corresponding integer sequences arise for a given irrational number $k$, an example being the Fibonacci numbers when one chooses $k=\phi$, the golden ratio. Here, we note that 414141 is a member of the integer sequence $\frac{1}{2}Q_{n}$, where $Q_{n}$ are the Pell-Lucas numbers. Just as successive ratios of the Fibonacci numbers form the best rational approximations for $\phi$ ($\lim_{n\rightarrow\infty}F_{n+1}/F_{n}=\phi$), similarly, $\lim_{n\rightarrow\infty}\frac{1}{2}Q_{n}/P_{n}=\sqrt{2}$, where $P_{n}$ here refers to the Pell numbers. Thus, if $L+1=\frac{1}{2}Q_{n}$ for some $n$, then $(L+1)\times\sqrt{2}\approx m\in\mathbb{Z}$, leading to the potential being nearly commensurate with the lattice. We confirm that for $k=\phi$, these spikes occur at system sizes corresponding to the Fibonacci numbers and other integers $m$ that happen to satisfy $m\times k\approx n$. This underlying commensurability is likely the culprit that leads to increased resonances and other strange behavior in the spectrum, leading to a spike in the quantities calculated here. This would be expected as a periodic system exhibits vastly different physics from an incommensurate one.
@@ III.4.1 LIOM Corrections: Anderson Model
Figure 3: Analysis of the Frobenius norm of the first-order corrections to the noninteracting LIOMs, $|\tilde{n}_{\alpha}^{(1)}|_{F}$, in the weakly interacting Anderson model. Each plot displays all orbitals $\alpha=1,\dots,L$ across the 500 disorder realizations used to generate the data shown in this figure. (a)-(d): Distributions of $\ln(\|\tilde{n}_{\alpha}^{(1)}\|_{F})$ at $W=0.6,1.6,2.6,3.1$ and for $\alpha=1,\dots,L$. Within each panel, each colored curve represents a different system size. We show calculations between $L=20$ and $L=100$ in steps of 101010. The inset in (d) shows the same data as in the main (d) panel, but with the vertical scale shown on log-scale. The red line in the inset represents a polynomial $\|\tilde{n}_{\alpha}^{(1)}\|_{F}^{-2}$ fit to the tail of the probability distribution of $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$— see the text surrounding Eq. (21) for details on the form plotted. The convergence of the distributions for large enough $W$ and $L$ indicates the presence of an extensive number of quasiconservation laws after interactions are added, even in the thermodynamic limit. (e): Median $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ as a function of $L$ on a log-log scale. Each colored curve represents a different disorder strength $W$, from $W=0.1$ to $W=3.1$ in steps of 0.50.50.5. One observes that the curves appear to saturate for large $L$, indicating that at least half of the LIOMs are conserved to second order when interactions are added. See Sec. III.4.1 for a full discussion of the results.
We present numerical calculations of the Frobenius norm of the LIOM corrections, $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$, in Fig. 3, for the weakly–interacting Anderson model. For each $L$, we sample 500500500 disorder realizations and record $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ for every orbital $\alpha=1,2,...,L$ in each disorder realization, giving a total of $500\times L$ datapoints in each distribution.
@@
Furthermore, in Fig. 4, we note that the $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ vs $L$ curves collapse when the x-axis is appropriately scaled by the participation ratio of the Anderson model, which scales approximately as $\mathrm{PR}(W)\propto W^{-1.78}$ in the range of $W$ considered here (see Appendix E for a more detailed explanation) and when the y-axis is scaled phenomenologically with $\sim\mathrm{PR}^{1.7}$. Here, the participation ratio is used as a proxy [^3] Some caution is needed in identifying the median participation ratio with the localization length in this analysis. First of all, both the participation ratio and the localization length depend on the energy of the noninteracting orbital, while our analysis does not consider the energy resolution of the LIOM corrections. Moreover, the power-law dependence of the localization length on the disorder strength is expected not to hold at large $W$, where a more refined ansatz is needed [75]. See Appendix E for another discussion. for the localization length $\xi$. The collapse of these curves suggests that at low disorders, the growth of $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ is a finite-size effect. If the system size $L$ exceeds the localization length, for any disorder strength $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ is expected to saturate to a finite value. Put together, these results suggest that even in the thermodynamic limit, a fraction of noninteracting LIOMs have finite and small corrections in the presence of interactions.
@@ III.4.2 LIOM Corrections: Quasiperiodic Model
Figure 5: Analysis of the Frobenius norm of the first-order corrections to the noninteracting LIOMs, $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$, for the weakly-interacting quasiperiodic model. (a)-(d): Distributions of $\ln(\|\tilde{n}_{\alpha}^{(1)}\|_{F})$ based on 500 phase shifts $\delta$, sampled at potential depths $h=0.6,1.6,2.6,3.1$, with $\alpha=1\dots L$. Each curve represents the distribution calculated at different system sizes, from $L=20$ to $L=100$ in steps of 10. The inset in (d) shows the same data as in the main (d) panel but with the vertical scale shown on a log scale. The red line in the inset represents a polynomial $\|\tilde{n}_{\alpha}^{(1)}\|_{F}^{-1.75}$ fit to the tail of the probability distribution of $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$— see the text surrounding Eq. (21) for details on the form plotted. Similarly to what was observed for the interacting Anderson model, the quasiperiodic model seems to admit an extensive number of quasiconservation laws for the localized phase $h>2$. (e): Median $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ as a function of $L$ on a log-log scale. Each colored curve represents a different potential depth $h$, from $h=0.1$ to $h=3.1$ in steps of 0.5. One observes that the curves for $h<2$ grow with $L$, indicating that the corrections are, unsurprisingly, delocalized. For $h>2$ the curves saturate, indicating that quasi-LIOMs persist when interactions are added to the localized regime of the quasiperiodic model. For a full discussion of the results, see Sec. III.4.2
@@ IV.1 Charge Transport Capacity
In this section, we introduce a measure of the charge transport capacity (for any process) across a link, inspired by the integrated energy current operator introduced in Ref. [76]. We consider an OBC chain and fix a site $i_{0}$ where $1\leq i_{0}\leq L-1$. Here, we fix $t=1$. We can define the charge current operator $\hat{J}_{i_{0}}$ across the link between $i_{0}$ and $i_{0}+1$, with rightward current being defined as positive:
@@
Continuing with general discussion, we can now more formally argue that in the MBL phase, assuming the LIOM description holds, the above $\hat{M}_{i_{0}}$ is related to LIOMs and noncommuting partners (raising/lowering operators) localized near $i_{0}$ and is localized near $i_{0}$, and hence the charge transport capacity is bounded by a fixed number in the thermodynamic limit. Indeed, let us take the LIOM description posited in Eq. (7) and express the total charge operator $\hat{N}_{\text{tot}}=\sum_{j}\hat{n}_{j}$ in terms of $\tilde{d}_{\alpha},\tilde{d}_{\alpha}^{\dagger}$. Here, $\tilde{d}_{\alpha},\tilde{d}_{\alpha}^{\dagger}$ are the annihilation and creation operators associated with the corresponding LIOM $\tilde{n}_{\alpha}$. Since $\hat{N}_{\text{tot}}$ commutes with the Hamiltonian, and assuming the Hamiltonian spectrum is non-degenerate, it follows that $\hat{N}_{\text{tot}}$ can be expressed as a sum of terms containing only $\tilde{n}_{\alpha}$’s. [^4] If we could in addition assume that there is an “adiabatic-type” one-to-one correspondence between the unperturbed LIOMs $\hat{d}_{\alpha}^{\dagger}\hat{d}_{\alpha}$ and the new LIOMs $\tilde{d}_{\alpha}^{\dagger}\tilde{d}_{\alpha}$, more specifically that there is a quasi-local unitary $U$ relating $\tilde{n}_{\alpha}\equiv\tilde{d}_{\alpha}^{\dagger}\tilde{d}_{\alpha}=U\hat{d}_{\alpha}^{\dagger}\hat{d}_{\alpha}U^{\dagger}$, requiring also that $U$ commutes with $\hat{N}_{\text{tot}}$, we would conclude $\hat{N}_{\text{tot}}=\sum_{\alpha}\tilde{n}_{\alpha}$. However, this specific form is not used in the argument that follows.
Consider now the parts $\hat{N}_{L}(i_{0})$ and $\hat{N}_{R}(i_{0})$, $\hat{N}_{L}(i_{0})+\hat{N}_{R}(i_{0})=\hat{N}_{\text{tot}}$. The assumed quasi-locality of the LIOMs implies that any local operator—in particular, each $\hat{n}_{j}$—is expressible in terms of quasi-local strings containing products of $\tilde{d}_{\alpha},\tilde{d}_{\alpha}^{\dagger},\tilde{n}_{\alpha}$ localized near $j$. For simplicity, let us assume that there is a maximal range $r_{\text{max}}$ for these strings. Now recall that $\hat{N}_{\text{tot}}$ contains only strings of $\tilde{n}$’s. Since any string in $\hat{N}_{R}(i_{0})$ that starts farther than $r_{\text{max}}$ away from $i_{0}$ is not present in $\hat{N}_{L}(i_{0})$ and hence must be the same as in $\hat{N}_{\text{tot}}$, such a string contains only $\tilde{n}$’s and hence commutes with $\hat{H}$. Hence $\hat{M}_{i_{0}}$, which is the off-diagonal part of $\hat{N}_{R}(i_{0})$, obtains contributions only from the strings in $\hat{N}_{R}(i_{0})$ that are localized near $i_{0}$, so $\hat{M}_{i_{0}}$ is indeed localized near $i_{0}$. See Fig. 7 for a visual depiction of this argument. We note that the restriction to strictly local strings (with size at most $r_{\text{max}}$) vs quasi-local (whose amplitudes decrease sufficiently fast with the string size) is not important for this argument.
@@
We note that the above argument relies on the validity of the LIOM model in the MBL phase. However, the possible double-logarithmic growth of the number entropy in the MBL regime, as reported in Ref. [77], suggests that more subtle transport phenomena may persist even deep within the localized phase; whether this is true is not entirely resolved in the literature [78, 79]. Regardless, we emphasize that our calculations of the charge transport in this paper are fully independent of any assumptions about the form of the LIOM description in the localized regime.